OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1)

OpenStudy (zehanz):

The general equation for a line through (a, b) and slope m is: y - b = m(x - a). Just fill in m = -2, a = 5 and b = -1 to see what eq you get!

OpenStudy (anonymous):

B. y – 1 = -2(x – 5)

OpenStudy (anonymous):

y - b = m(x - a). b = -1 m = -2 a = 5 y + 1 = -2(x -5) Which is the 1st one.

OpenStudy (anonymous):

Determine if lines that pass through the given points are parallel, perpendicular or neither. Line A: (5, 4) and (7, 5) Line B: (4, -5) and (8, -3) A. Parallel B. Perpendicular C. Neither D. Cannot be determined

OpenStudy (anonymous):

Find the slope for each line: A and B Line A: 5 - 4/7 - 5 = 1/2 Line B: -3 + 5/ 8 - 4 = 2/4 = 1/2 Slope of Line A = Slope of Line B They are PARALLEL.

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (zehanz):

About the parallel/perp lines: 1. calc the slope by delta y / delta x: Line A: (4-5)/(5-7)= -1/-2 = -1/2 Line B: (-5--3)/(4-8)=-2/-4= 1/2 Two lines are parallel if they have the same slope --> NOT Two lines are perp. if the product of their slopes = -1 --> NOT They have different slopes and are not perp, so C. neither

OpenStudy (zehanz):

Oops, Line A has also slope 1/2, so they are A. Parallel...

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (zehanz):

Next one: parallel, so same slope = 3. Equation is now y = 3x + b. Find b this way: b = y - 3x. Fill in the values for x and y from the given point (5,2), and you will know the right answer!

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (zehanz):

b = 2 - 3*5 = 2 -15 = -13, so the answer is B.

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). When you say parallel, the slope must be equal. In this case m = 3 for both lines. We now need the x and y intercepts. Recall that the slope intercept form is y = mx + c We have m = 3 So the equation is now y = 3x + c We have the values for x and y x = 5; y = 2 Substitute them in the equation above: 2 = 3(5) + c 2 = 15 + c 2 - 15 = c -13 = c Now input the final equation: y = 3x - 13 Which is B.